OpenStudy (anonymous):

Simplify each product. Write in standard form. 18. 6g^2 (g - 8) = 20. -2n^2 (5n - 9 + 4n^2) =

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

u know the way to solve it?

OpenStudy (anonymous):

@ArifL Not really. I need someone to talk me threw it.

OpenStudy (anonymous):

just multiply the outer number to each one inside

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@ArifL Can you show me please so I can understand?

OpenStudy (anonymous):

@carsonce

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

@carsonce Can you help me please?

OpenStudy (anonymous):

\[6g ^{2}(g-8) \]\[(6g ^{2}\times g)+(6g ^{2}\times-8)\]

OpenStudy (anonymous):

There you go ^

OpenStudy (anonymous):

@ArifL Can you explain to me how you got that answer? And is (6g^2 times g) + (6g^2 times -8) the answer?

OpenStudy (anonymous):

no, it not the answer

OpenStudy (anonymous):

it the first step

OpenStudy (anonymous):



OpenStudy (anonymous):

So what is the next step?

OpenStudy (anonymous):

just finish it

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ArifL 12g^2 + (-8) is that correct?

OpenStudy (anonymous):

yeah, sorry im not that good at teaching

OpenStudy (anonymous):

no it 6g^3-48g^2

OpenStudy (anonymous):

Can you explain to me how that is the answer?

OpenStudy (anonymous):

How is it 6g^3 ? How does the exponent change?

OpenStudy (anonymous):

\[a ^x \times a^y= a ^{x+y}\]

OpenStudy (anonymous):

it just count if it is the same variable

OpenStudy (anonymous):

@ArifL If you're adding the exponents shouldn't it be 6g^4 ?

OpenStudy (anonymous):

\[6g^2\times g^1\]

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @ArifL \[a ^x \times a^y= a ^{x+y}\] \(\color{blue}{\text{End of Quote}}\) look at the exponent rule provided by arifl when we multiply same bases we should `add` their exponents

OpenStudy (anonymous):

6g^2 (g - 8) = (6g^2 times g) + (6g^2 times -8) = 6g^3 - 48g^2 is that in standard form?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

just that is good

OpenStudy (anonymous):

how would i explain it, u just need to multiply it

OpenStudy (anonymous):

@MrNood can u help us?

OpenStudy (anonymous):

@Nnesha can u help us again?

Nnesha (nnesha):

yes ?

Nnesha (nnesha):

it's distributive property multiply each term in the parentheses by outside ter m

OpenStudy (anonymous):

@Nnesha THANK YOU!!

Nnesha (nnesha):

hmm yw??

Nnesha (nnesha):

 just like arif mentioned multiply each term by outside term :=))